OpenStudy (anonymous):

What is the x-coordinate of the solution to the system of equations? 3x – 2y = –30 6x + 3y = 3 _________

OpenStudy (mismycatmin):

1) 3x − 6y = 3 . . 7x − 5y = −11 -5(3x − 6y = 3) 6(7x − 5y = −11) -15x + 30y = -15 42x − 30y = −66 -------------------------- 27x = -81 x = -3 If x = -3 and 7x − 5y = −11, then 7x − 5y = −11 7(-3) − 5y = −11 -21 - 5y = -11 -5y = 10 y = -2 <==ANSWER for y (-3,-2) is the solution

OpenStudy (mismycatmin):

2) 3x + y = 11 . –3x + 2y = 1 ---------------------- 3y = 12 y = 4 <==ANSWER (7/3,4) is solution

OpenStudy (mismycatmin):

3) The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations auditioned today? x + y = 1150 3x - 2y = 1300 2( x + y = 1150) 3x - 2y = 1300 2x + 2y = 2300 3x - 2y = 1300 ----------------------- 5x = 3600 x = 720 <==ANSWER for one step equations

OpenStudy (mismycatmin):

4) 3x + y = 6 . . 6x + 2y = 8 -2(3x + y = 6) . .6x + 2y = 8 -6x - 2y = -12 6x + 2y = 8 -------------------- 0 = -4 That's never true, so there is no x value as there is no solution to the system. <==ANSWER

OpenStudy (mismycatmin):

5) 4x + 2y = 6 . . x – y = 3 4x + 2y = 6 2(x – y = 3) 4x + 2y = 6 2x – 2y = 6 ------------------ 6x = 12 x = 2 <==ANSWER

OpenStudy (mismycatmin):

hope that is right

OpenStudy (anonymous):

So x=2 not 3?

OpenStudy (mismycatmin):

and thanks for the medal btw i was loveheart but i got hacked out of my account and i have had this account so its Kimberly thanks for trying to get him out

OpenStudy (mismycatmin):

and yes x = 2

OpenStudy (anonymous):

Oh, hey kimberly!

OpenStudy (mismycatmin):

hey i am not shear if you got me messages from arlyar when i do get that account back i will say

OpenStudy (mismycatmin):

if you do not bolev it is me i will change i pic to the one that i had in the first place

OpenStudy (mismycatmin):

@NormaValenzuela are you still there

OpenStudy (anonymous):

I believe its you.

OpenStudy (anonymous):

and thxs for the help...but i got more questions.

OpenStudy (mismycatmin):

ok close this one and tag me